OpenStudy (anonymous):

Find x-component of vector v⃗ =(250m/s , 30∘ above + x-axis)

OpenStudy (theeric):

Hi! Have you used trigonometric functions before, like cosine? Cosine is usually written as "\(cos\)" and is a function of an angle, usually \(\theta\). So you might recognize "\(cos\left(\theta\right)\)".

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so yes x = cos, y = sin i'm asked to find x component and y component of vector is it as simple as cos30(250) and sin30(250)? i'm just thrown off because 250 is a velocity

OpenStudy (theeric):

Haha, I got ya. The velocity has magnitude and direction, so it's exactly like any other vector. Just different units from things like force, or work. So you will do \(250\ [m/s] \times cos\left( 30^\circ\right)\) for the \(x\)-component.

OpenStudy (anonymous):

x=216.51 y=125

OpenStudy (theeric):

I agree! \(\huge\color{blue}{\large\ \ o\ o\ \\\smile}\)

OpenStudy (anonymous):

thanks do you have time for one more

OpenStudy (theeric):

Possibly. Post it as another question so other can see it to help out too!

OpenStudy (anonymous):

Find x-component of vector a⃗ =(6.0m/s2 , - y-direction) this is all the problem gives me. i should i approach the problem.

OpenStudy (theeric):

I see.

OpenStudy (anonymous):

negative y

OpenStudy (anonymous):

its looking for a sub y

OpenStudy (theeric):

Can you draw a picture of the vector? It might help make sense of it visually. Otherwise, you could just realize that if the entire vector is only in the \(y\) or negative \(y\) direction, then there is no \(x\) component.

OpenStudy (anonymous):

so its downward vector going down the y axis

OpenStudy (anonymous):



OpenStudy (theeric):

Yup! Right on the \(y\)-axis. So, there is no \(x\)-component at all.

OpenStudy (theeric):

You can also do this:\(\overrightarrow a\)is at an angle of \(270^\circ\). What is the \(x\)-component, now? \(\overrightarrow{a_x} = a\ cos\left( 270^\circ\right) \stackrel{?}{=}\)